# Perception, Reason, Think, and Plan: A Survey on Large Multimodal Reasoning Models

### v1

@@ 3.3.2 Multimodal-O1
With the rise of OpenAI o1, which sparked widespread interest in large reasoning models, open-source reproductions such as Marco-o1 (Zhao et al., 2024c) and llamaberry (Zhang et al., 2024b) utilizing CoT fine-tuning began to emerge. CoT fine-tuning activates the model’s inherent slow thinking ability through training methods. Compared to traditional CoT approaches, it enhances the model’s reasoning capabilities on open-ended questions, introducing mechanisms for self-reflection and error correction. LLaVA-CoT (Xu et al., 2024b), LlamaV-o1 (Thawakar et al., 2025), RedStar (Xu et al., 2025a) and Mulberry (Yao et al., 2024a) extend the reasoning paradigm to the multimodal domain. In contrast to the two-stage reasoning paradigm of ’Thinking -> Answer’ in text domains, these works expand the reasoning process to a four-stage approach includes Summary (Rationale), Caption, Thinking and Answer.
Building on CoT fine-tuning, testing-time scaling with various reasoning strategies is also an important method to enhance reasoning capabilities. Best-of-N sampling generates multiple responses for a given prompt, expanding the search space to identify better solutions. Beam Search, on the other hand, does not generate a complete response in one pass but instead selects the most promising intermediate outputs at each step using scoring. LLaVA-CoT (Xu et al., 2024b) and LlamaV-o1 (Thawakar et al., 2025) apply this method to strengthen reasoning abilities. Monte Carlo Tree Search (MCTS) allows for parallel exploration of multiple solution paths, ensuring a more refined search process compared to Beam Search. Marco-o1 (Zhao et al., 2024c), llamaberry (Zhang et al., 2024b) and Mulberry (Yao et al., 2024a) have successfully integrated this approach into the generation process of reasoning models.
@@
Table 6: Apporach of Multimodal-R1. It mainly employs reinforcement learning approaches to improve the reasoning capability of large multimodal models.
@@
InternVL2.5 (2024g) InternVL MMPR (3m) MPO (DPO) T, I VQA - - (unknown)
@@
VLM-R1 (2025) Qwen2.5-VL coco , LISA , Refcoco GRPO T, I Grounding ,Math , Open-Vocabulary Detection trl No Rule-base
R1-V (2025b) Qwen2-VL CLEVR , GEOQA GRPO T, I Counting , Math trl No Rule-base
@@
Vision-R1 (2025b) Qwen2.5-VL LLaVA-CoT , Mulberry GRPO T, I Math - Yes Rule-base
Visual-RFT (2025f) Qwen2-VL coco , LISA , … GRPO T, I Detection , Classification trl No Rule-base
R1-OneVision (2025e) Qwen2.5-VL R1-Onevision-Dataset GRPO T, I Math , Science , General , Doc - Yes Rule-base
@@
VisualThinker-R1-Zero (2025) Qwen2-VL SAT dataset GRPO T, I Spatial Reasoning trl No Rule-base
R1-Omni (2025c) HumanOmni MAFW , DFEW GRPO T, I, A, V emotion recognition trl Yes Rule-base
OThink-MR1 (2025e) Qwen2.5-VL CLEVR , GEOQA GRPO T, I Counting , Math - No Rule-base
@@
Curr-ReFT (2025) Qwen2.5-VL RefCOCOg , Math360K , Geo170K GRPO T,I Detection , Classification , Math Curr-RL No RM
@@
VisRL (2025f) Qwen2.5-VL VisCoT DPO T,I VQA trl Yes RM
@@
R1-VL (2025c) Qwen2-VL Mulberry-260k StepGRPO T,I Math , ChartQA not release No Rule-base

### v2

@@ 3.3.2 Multimodal-O1
With the rise of OpenAI o1, which sparked widespread interest in large reasoning models, open-source reproductions such as Marco-o1 (Zhao et al., 2024c) and llamaberry (Zhang et al., 2024b) utilizing CoT fine-tuning began to emerge. Compared to traditional CoT approaches, CoT fine-tuning enhances the model’s reasoning capabilities on open-ended questions by incorporating mechanisms for self-reflection and error correction. LLaVA-CoT (Xu et al., 2024b), LlamaV-o1 (Thawakar et al., 2025), RedStar (Xu et al., 2025a) and Mulberry (Yao et al., 2024a) extend the reasoning paradigm to the multimodal domain. In contrast to the two-stage reasoning paradigm of ’Thinking -> Answer’ in text domains, these works expand the reasoning process to a four-stage approach that includes Summary (Rationale), Caption, Thinking and Answer.
Building on CoT fine-tuning and testing-time scaling with various reasoning strategies is also an important method to enhance reasoning capabilities. Best-of-N sampling generates multiple responses for a given prompt, expanding the search space to identify better solutions. Beam Search, on the other hand, does not generate a complete response in one pass but instead selects the most promising intermediate outputs at each step using scoring. LLaVA-CoT (Xu et al., 2024b) and LlamaV-o1 (Thawakar et al., 2025) apply this method to strengthen reasoning abilities. Monte Carlo Tree Search (MCTS) allows for parallel exploration of multiple solution paths, enabling more comprehensive exploration and higher-quality solutions compared to Beam Search. Marco-o1 (Zhao et al., 2024c), llamaberry (Zhang et al., 2024b), and Mulberry (Yao et al., 2024a) have successfully integrated this approach into the generation process of reasoning models.
@@
Table 6: Approaches of Multimodal-R1. It mainly employs reinforcement learning approaches to improve the general reasoning capability of large multimodal models.
@@
InternVL2.5 (2024g) InternVL MMPR (3M) MPO (DPO) T, I VQA - - (unknown)
@@
VLM-R1 (2025b) Qwen2.5-VL coco , LISA , Refcoco GRPO T, I Grounding ,Math , Open-Vocabulary Detection trl No Rule-base
R1-V (2025c) Qwen2-VL CLEVR , GEOQA GRPO T, I Counting , Math trl No Rule-base
@@
Vision-R1 (2025c) Qwen2.5-VL LLaVA-CoT , Mulberry GRPO T, I Math - Yes Rule-base
Visual-RFT (2025l) Qwen2-VL coco , LISA , … GRPO T, I Detection , Classification trl No Rule-base
R1-OneVision (2025g) Qwen2.5-VL R1-Onevision-Dataset GRPO T, I Math , Science , General , Doc - Yes Rule-base
@@
VisualThinker-R1-Zero (2025b) Qwen2-VL SAT dataset GRPO T, I Spatial Reasoning trl No Rule-base
STAR-R1 (2025q) Qwen2.5-VL-7B TRANCE (13.5k) GRPO T, I Spatial Reasoning vLLM No Rule-base
R1-Omni (2025d) HumanOmni MAFW , DFEW GRPO T, I, A, V emotion recognition trl Yes Rule-base
OThink-MR1 (2025k) Qwen2.5-VL CLEVR , GEOQA GRPO T, I Counting , Math - No Rule-base
@@
Reason-RFT (2025) Qwen2.5-VL CLEVR-Math, Super-CLEVR, GeoMath, Geometry3K, TRANCE GRPO T, I Counting, Structure Perception, Spatial Transformation trl No Rule-base
VL-Rethinker (2025i) Qwen2.5-VL MathVista, MathVerse, MathVision, MMMU-Pro, EMMA, MEGA GRPO+SSR T, I Mathematical, Scientific, Real-world Reasoning trl No Rule-base
Curr-ReFT (2025b) Qwen2.5-VL RefCOCOg , Math360K , Geo170K GRPO T,I Detection , Classification , Math Curr-RL No RM
@@
VisRL (2025l) Qwen2.5-VL VisCoT DPO T,I VQA trl Yes RM
@@
R1-VL (2025g) Qwen2-VL Mulberry-260k StepGRPO T,I Math , ChartQA not release No Rule-base
SARI (2025) Qwen2-Audio-7B-Instruct MMAU (32k) GRPO A, T MCQA not release No Rule-based
SpaceR (2025) Qwen2.5-VL-7B SpaceR-151k SG-RLVR(GRPO) T,I,V Video Spatial Reasoning - Yes Rule-base
Table 7: Approaches of Multimodal-R1 on image understanding and generation. It mainly employs reinforcement learning approaches to improve the generative reasoning of large multimodal models.
Cold ViCrit (2025s) Qwen2.5-VL-3B/72B-Instruct PixMo-Cap GRPO T, I Hallucination Detection not release No Rule-base
Vision Matters (2025o) Qwen2.5-VL-Instruct Geometry3K, TQA, GeoQA, Math8K, M3CoT GRPO + DPO T, I Math MS-Swift, EasyR1 No RM
ViGaL (2025b) Qwen2.5-VL-7B-Instruct Sampled from game: Snake(36K), Rotation(36K) RLOOT T, I Visual Games OpenRLHF No Rule-base
RAP (2025f) Qwen2.5-VL-3B, Qwen2.5-VL-7B MM-Eureka GRPO, RLOOT T, I Data Selection EasyR1 No Not metion
RACRO (2025b) Qwen2.5-VL(3B, 7B, 32B) ViRL39K CROT T, I change reasoner without re-alignment verl No combine
Revisual-R1 (2025e) Qwen2.5-VL-7B-Instruct GRAMMAR GRPO T, I Math EasyR1 Yes Rule-base
Rex-Thinker (2025c) Qwen2.5-VL-7B HumanRef-CoT GRPO T, I Object Referring (REC) verl Yes RM
ControlThinker (2025) ControlARC COCOStuff, MultiGen-20M GRPO T, I Image Editing not release Yes RM
SynthRL (2025g) Qwen2.5-VL-7B-Instruct MMK12, A-MMK12 GRPO T, I Math verl No RM
SRPO (2025) Qwen-2.5-VL-7B, Qwen-2.5-VL-32B Mulberry dataset (260K), MathV360K, LLaVA-CoT (100K) , ScienceQA , etc. GRPO T, I Math verl Yes RM
ReasonGen-R1 (2025p) Janus-Pro-7B LAION-5B GRPO T, I Text to Image Generation verl Yes RM
MoDoMoDo (2025c) Qwen2-VL-2B-Instruct COCO, LISA, GeoQAV, SAT, ScienceQA GRPO T, I General Visual Reasoning trl No RM
DINO-R1 (2025a) MM-Grounding-DINO Objects365 GRPO T, I Object Detection not release Yes RM
VisualSphinx (2025c) Qwen2.5-VL-7B VISUALSPHINX GRPO T, I visual logic puzzle, math verl No Rule-base
PixelThink (2025q) Qwen2.5-VL-7B, SAM2-Large RefCOCOg GRPO T, I Segmentation verl No Rule-base
ViGoRL (2025) Qwen2.5-VL-3B, Qwen2.5-VL-7B SAT-2, OS-ATLAS, ICAL, Segment Anything GRPO T, I spatial reasoning, web grounding, web action, etc. verl Yes Rule-base
Jigsaw-R1 (2025y) Qwen2.5-VL-7B, Qwen2.5-VL-3B, Qwen2-VL-2B, InternVL2.5-2B COCO, CV-Bench, MMVP, SAT, Super-CLEVR GRPO T, I jigsaw puzzles trl No Rule-base
UniRL (2025) Show-o, Janus COCO, GPT4o-Generated GRPO T, I Image Understanding & Generation not release Yes Rule-base
cadrille (2025) Qwen2-VL-2B DeepCAD DPO, GRPO T, I Computer-Aided Design not release Yes Rule-base
MM-UPT (2025d) Qwen2.5-VL-7B Geo3K, GeoQA, MMR1 GRPO T, I Math verl No Rule-base
RL-with-Cold-Start (2025d) Qwen2.5-VL-3B, Qwen2.5-VL-7B Geometry3K, GeoQA, GeoQA-Plus, Geos, AI2D, etc. GRPO T, I Multimodal Reasoning, Math verl Yes Rule-base
VRAG-RL (2025p) Qwen2.5-VL-3B, Qwen2.5-VL-7B ViDoSeek, SlideVQA, MMLongBench GRPO T, I Visually Rich Information Understanding verl Yes RM + Rule-base
MLRM-Halu (2025a) Qwen2.5-VL(3B,7B) MMMU, MMVP, MMBench, MMStar, etc. GRPO T, I reasoning, perception not release Yes Rule-base
Active-O3 (2025d) Qwen2.5-VL-7B SODA, LVIS GRPO T, I active perception not release Yes RM
RLRF (2025) Qwen2.5-VL(3B,72B), Qwen3-8B SVG-Stack GRPO T, I Inverse rendering not release Yes RM
VisTA (2025e) Qwen2.5-VL-7B ChartQA, Geometry3K GRPO T, I Visual Reasoning, Tool Selection OpenR1 Yes RM+Rule-base
Table 8: Approaches of Multimodal-R1 on image-text reasoning and image generation. It mainly employs reinforcement learning approaches to improve the reasoning capability of large multimodal models.
SATORI-R1 (2025a) Qwen2.5-VL-Instruct-3B Text-Total, ICDAR2013, ICDAR2015, etc. GRPO T, I task-critical regions, answer accuracy not release No RM
URSA (2025b) Qwen2.5 Math-Instruct, SAM-B+SigLIP-L DualMath-1.1M GRPO T, I data reasoning, reward hacking URSA No RM
v1 (2025) Qwen2-VL(7B,72B), Qwen2.5-VL(7B,72B) v1g NoT I retrieve regions None No No
GRE Suite (2025d) Qwen2.5VL(3B,7B,32B) Im2GPS3k, GWS15k GRPO T, I reasoning location LLaMA-Factory Yes RM+Rule-base
V-Triune (2025c) Qwen2.5-VL-7B/32B-Instruct mm_math, geometry3k, mmk12, PuzzleVQA, etc. GRPO T, I intensive perception verl Yes RM
RePrompt (2025c) Qwen2.5 7B GenEva GRPO T, I image generation trl Yes RM
GoT-R1 (2025a) Qwen2.5VL-7B JourneyDB-GoT, FLUX-GoT GRPO T, I semantic-spatial reasoning not release No RM
SophiaVL-R1 (2025a) Qwen2.5-VL-7B-Instruct SophiaVL-R1-130k GRPO T, I reasoning-specific, general understanding VeRL No RM+Rule-base
R1-ShareVL (2025a) Qwen2.5-VL-7B/32B MM-Eureka GRPO T, I General Visual Reasoning EasyR1 No Rule-base
VLM-R3 (2025a) Qwen2.5-VL-7B VLIR GRPO T, I Region Recognition and Reasoning DeepSpeed Yes Rule-base
TON (2025l) Qwen-2.5-VL-Instruct-3B/7B CLEVR, Super-CLEVR, GeoQA, AITZ GRPO T, I counting, navigation, math reasoning vLLM Yes Rule-base
Pixel Reasoner (2025a) Qwen2.5-VL-7B SA1B, FineWeb and STARQA GRPO T, I pixel-space reasoning OpenRLHF No Rule-base
GRIT (2025b) Qwen2.5-VL-3B, InternVL-3-2B VSR, TallyQA, GQA, MME, MathVista, etc. GRPO T, I visual grounding, multi-step reasoning Deepspeed Zero2 No Rule-base+RM
VARD (2025a) SD 1.4 SCOPe, Pick-a-Pic, ImageRewardDB DDPO w/o KL I image generation not release No RM
Chain-of-Focus (2025m) Qwen2.5-VL-7B MM-CoF, SA_1B, TextVQA, m3cot, V, POPE GRPO T, I visual search and reasoning not release Yes Rule-base
Visionary-R1 (2025) Qwen2.5-VL-3B A-OKVQA, ChartQA, AI2D, ScienceQA, etc. GRPO T, I VQA not release No Rule-base
VisualQuality-R1 (2025f) Qwen2.5-VL-7B KADID-10K, SPAQ GRPO T, I image quality scoring not release No Rule-base
DeepEyes (2025f) Qwen2.5-VL-7B V* training set, ArxivQA, ThinkLite-VL GRPO T, I Multimodal Reasoning verl No Rule-base
Visual-ARFT (2025m) Qwen2.5-VL(3B,7B) MAT-Search, MAT-Coding, 2WikiMultihopQA, etc. GRPO T, I Multimodal Agentic Reasoning not release No Rule-base
UniVG-R1 (2025) Qwen2-VL-2B 7B MGrounding-630k, RefCOCO/+/g, etc. GRPO T, V, I Visual Grounding (Multi-image, Complex) Open-R1 Yes RM+Rule-base
G1 (2025b) Qwen2.5-VL-7B 128 parallel games, 500 training steps GRPO T, I Interactive Game Decision-Making EasyR1 Yes Rule-base
VisionReasoner (2025j) Qwen2-VL COCO, RefCOCO(+/g), ReasonSeg, etc. GRPO T, I detection, segmentation, counting not release No Rule-base
VPRL (2025g) LVM-3B FrozenLake, Maze, MiniBehavior GRPO I, V Visual Spatial Planning not release Yes Rule-base
GuardReasoner-VL (2025h) Qwen2.5-VL Instruct 3B/7B GuardReasoner-VLTrain GRPO T, I Moderation (Harmfulness Detection) EasyR1 Yes Rule-base
Table 9: Approaches of Multimodal-R1 on mathematic, code, and multimodal grounding. It mainly employs reinforcement learning approaches to improve the reasoning capability of large multimodal models.
OpenThinkIMG (2025b) Qwen2-VL-2B-Instruct CHARTGEMMA GRPO T, I Chart Reasoning Open-R1 Yes Rule-base
DanceGRPO (2025) Stable Diffusion, HunyuanVideo, FLUX, etc. curated prompt dataset, VidProM GRPO T, I T2V, I2V, T2I Generation fastvideo No RM
Flow-GRPO (2025b) SD3.5-M GenEval, OCR, from pickscore GRPO T, I Composition Image Gen, Text Rendering, etc. no release No RM, Rule-base
X-Reasoner (2025e) Qwen2.5-VL-7B-Instruct OpenThoughts, Orz-math, MedQA GRPO T, I Generalization across domains no release No Rule-base
T2I-R1 (2025b) Janus-Pro-7B T2I-CompBench GRPO T, I Text-to-Image Generation Open-R1 No RM
RM-R1  (2025i) Qwen-Instruct (7B/14B/32B), DeepSeek-Distilled-Qwen (7B/14B/32B) Skywork-Reward-Preference, Code-Preference-Pairs, Math-DPO-10K GRPO T Reward Modeling verl Yes RM
AutoThink (2025c) DeepSeek-R1-Distill-Qwen-1.5B MATH, Minerva, Olympiad, AIME24, AMC23 GRPO T Mathematical Reasoning verl No RM
WEBAGENT-R1 (2025f) Qwen2.5-3B/Llama3.1-8B WebArena-Lite M-GRPO T web tasks no release Yes RM
R1-Code-Interpreter (2025k) Qwen-2.5-(3B, 7B, 14B) SymBench, BIG-Bench-Hard, Reasoning-Gym GRPO T planning verl Yes RM
Mixed-R1 (2025d) Qwen2.5-VL-(3B, 7B) Mixed-45K GRPO T, I, V reasoning no release Yes RM + Rule-base
Router-R1  (2025e) Qwen2.5-3B-Instruct , LLaMA-3.2-3B-Instruct Natural Questions, TriviaQA, PopQA; HotpotQA, 2WikiMultiHopQA, Musique, Bamboogle PPO T Multi-hop Question Answering verl Yes RM + Rule-base
FinLMM-R1 (2025) Qwen2.5-VL-3B FinData GRPO T, I Reasoning TAR-LMM No RM
GVM-RAFT (2025b) Qwen2.5-Math-1.5B and Qwen2.5-Math-7B Numina-Math Dynamic RAFT T Math verl No Rule-base
LoVeC (2025b) Llama-3-8B-Instruct and Gemma-2-9B-It WildHallu, Bios, PopQA GRPO, DPO, and ORPO T long-form generation TRL/vLLM No Rule-base+RM
Critique-GRPO (2025j) Qwen2.5-7B-Base and Qwen3-8B-Base OpenR1-Math-220k GRPO T mathematical, STEM, and general reasoning verl Yes RM
ReCode (2025b) Qwen-2.5-Coder-7B-Instruct and DeepSeekv1.5-Coder-7B-Instruct construct own training dataset GRPO, DAPO T code generation not release No Rule-base
ShapeLLM-omni (2025a) Qwen-2.5-VL-Instruct-7B 3D-Alpaca Not explicitly mentioned (Uses autoregressive models) T, I, 3D 3D Generation, 3D Understanding, 3D Editing not release No Rule-based
Omni-Perception (2025x) PD-RiskNet Simulated with LiDAR simulation toolkit PPO LiDAR + Proprioception Locomotion + Collision Avoidance not release No RM
Omni TM-AE (2025) TM-AE One Billion Word Benchmark, IMDb, News20 Tsetlin Automaton Logic T Similarity, Classification, Clustering no release No RM
DeepVideo-R1 (2025) Qwen2.5-VL-Instruct SEED-Bench-R1-Train + NExTGQA GRPO T, I, V Video VQA not release No Rule-base
ReFoCUS (2025) LLaVA-OV / InternVL ReFoCUS-962K GRPO T, I, V Video VQA not release No RM
Table 10: Approaches of Multimodal-R1 on video, medical and audio. It mainly employs reinforcement learning approaches to improve the reasoning capability of large multimodal models on a specific modality.
TW-GRPO (2025) Qwen2.5-VL-Instruct CLEVRER dataset GRPO T, I, V Video VQA trl No Rule-base
Spatial-MLLM (2025a) Qwen2.5-VL-Instruct Spatial-MLLM-120k GRPO T, I, V Spatial not release Yes Rule-base
VAU-R1 (2025c) Qwen2.5-VL-Instruct VAU-Bench-Train GRPO T, I, V Video Grounding trl No Rule-base
MUSEG (2025a) Qwen2.5-VL-Instruct E.T. Instruct 164k + CharadesSTA GRPO T, I, V Video VQA + Video Grounding trl No Rule-base
VerIPO (2025m) Qwen2.5-VL-Instruct DAPO-Math + ViRL39K + VQA-Video-24K GRPO + DPO T, I, V Video VQA + Spatial OpenRLHF No Rule-base
SpaceR (2025) Qwen2.5-VL-Instruct SpaceR-151k GRPO T, I, V Spatial + VideoVQA trl No Rule-base
TinyLLaVA-Video-R1 (2025k) Qwen2.5-VL-Instruct NextQA GRPO T, I, V Video VQA trl Yes Rule-base
VideoChat-R1 (2025j) Qwen2.5-VL-Instruct Charade - STA + NExTGQA + FIBER-1k + VidTAB GRPO T, I, V Video Grounding + Video VQA trl No Rule-base
R1-Zero-VSI (2025) Qwen2-VL-Instruct VSI-100k GRPO T, I, V Spatial not release No Rule-base
GRPO-CARE (2025j) Qwen2.5-VL-Instruct SEED-Bench-R1-Train GRPO T, I, V Video VQA + Spatial trl No Rule-base
Time-R1 (2025t) Qwen2.5-VL-Instruct Temporal-Spatial Video Datasets GRPO T, I, V Video Grounding trl Yes Rule-base
Lingshu (2025e) Qwen2.5-VL-Instruct 3.75M open-source medical samples and 1.30M synthetic medical samples / MedEvalKit GRPO T, I multimodal QA, medical report generation not release Yes Rule-base
Patho-R1 (2025h) OpenAI-CLIP/Qwen2.5VL PubMed, Quilt, PathGen GRPO+DAPO T, I Open-ended/Close-ended VQA verl Yes Rule-base
ChestX-Reasoner (2025c) Qwen2VL-7B Medical Datasets GRPO T, I single/binary disease diagnosis verl Yes Rule-base
Med-R1 (2025) Qwen2-VL-2B-Instruct OmniMedVQA GRPO T, I medical VQA not release Yes Rule-base
MedVLM-R1 (2025b) Qwen2-VL-2B HuatuoGPT-Vision GRPO T, I Radiological VQA not release Yes Rule-base
SARI (2025) Qwen2-Audio-7B-Instruct/ Qwen2.5-Omni AudioSet, MusicBench, Meld, AVQA GRPO T, A Audio QA trl No Rule-base
R1-AQA (2025d) Qwen2-Audio-7B-Instruct AVQA GRPO T, A Audio QA trl Yes Rule-base
Audio-Reasoner (2025c) Qwen2-Audio-7B-Instruct AVQA GRPO T, A Audio QA not release Yes Rule-base
Omni-R1 (2025) Qwen2.5-Omni AVQA GRPO T, A Audio QA not release No Rule-base
WavReward (2025) Qwen2.5-Omni-7B-Think ChatReward-30K PPO T, A end-to-end dialogue not release No Rule-base
SoundMind (2025) Qwen2.5-Omni-7B Audio Logical Reasoning (ALR) REINFORCE++ T, A Audio text bimodal reasoning verl No Rule-base
AudSemThinker (2025) Qwen2.5-Omni-7B AUDSEM GRPO T, A semantic audio reasoning trl No Rule-base
AV-Reasoner (2025b) Ola-Omni7B AVQA, Music AVQA, AVE, UnAV, LLP, AVSS-ARIG, DVD-Counting, RepCount GRPO T, I, V, A Counting, Video VQA, Reasoning trl Yes Rule-base
Omni-R1 (ZJU) (2025a) Qwen2.5-Omni-7B RefAVS, ReVOS, MeViS, refCOCOg GRPO T, V, A Multimodal Segmentation (AVS, VOS) trl Yes Rule-base
Omni-R1 (MIT) (2025) Qwen2.5-Omni-7B AVQA-GPT, VGGS-GPT GRPO T, A Audio QA not release no RM
EchoInk-R1 (2025b) Qwen2.5-Omni-7B AVQA-R1-6K GRPO T, I, A Audio VQA trl no Rule-base
Skywork-VL Reward (2025r) Qwen2.5-VL-7B-Instruct LLaVA-Critic-113k, Skywork-Reward-Preference-80Kv0.2, RLAIF-V-Dataset MPO T, I VQA, Math, Science, Reasoning not release no Rule-base
Table 11: Approaches of Multimodal-R1 on GUI (agentic) and reward models. It mainly employs reinforcement learning approaches to improve the interactive reasoning capability of large multimodal models.
UnifiedReward-Think (2025u) UnifiedReward HPD(25.6K), EvalMuse(3K), OpenAI-4o_t2i_human_preference (6.7K), VideoDPO (10K), Text2Video Human Preference (5.7K), ShareGPTVideo-DPO (17K) GRPO T, I, V Video/Image Understanding, Reward Assessment trl yes Rule-base
R1-Reward (2025n) QwenVL-2.5-7B-Instruct RLAIF-V, VL-Feedback, POVID, WildVision-Battle StableReinforce (Reinforce++ variant) T, I, V Video/Image Understanding, Reward Assessment OpenRLHF yes Rule-base
ComfyUI-R1 (2025j) Qwen2.5-Coder-7B-Instruct no release GRPO T, I, V workflow generation verl yes Rule-base
GUI-Critic-R1 (2025) Qwen2.5-VL-7B-Instruct GUI-Critic-Train GRPO T, I GUI Operation Error Detection and Correction no release yes Rule-base
ARPO (2025a) UI-Tars-1.5-7B OS World GRPO T, I GUI automation verl no Rule-base
GUI-G1 (2025e) Qwen2.5-VL-3B-Instruct UI-BERT, OS-Atlas (17K) GRPO T, I GUI grounding no release no Rule-base
UIShift (2025c) Qwen2.5-VL-3/7B-Instruct, no release GRPO T, I GUI automation and grounding VLM-R1 no Rule-base
MobileIPL (2025b) Qwen2-VL-7B MobileIPL-dataset DPO T, I GUI automation no release yes Rule-base
InfiGUI-R1 (2025i) Qwen 2.5-VL-3B-Instruct AndroidControl, ScreenSpot , ScreenSpot-Pro, Widget-Caption, COCO RLOO T, I GUI automation, GUI grounding no release no Rule-base
GUI-R1 (2025c) QwenVL 2.5-3B/7B GUI-R1-3K GRPO T, I GUI automation, GUI grounding EasyR1 no Rule-base
UI-R1 (2025c) Qwen2.5-VL-3B ScreenSpot (mobile subset), AndroidControl (1K) GRPO T, I GUI Action Prediction, GUI grounding no release no Rule-base